### v2

@@ Abstract
We explore the application of machine learning algorithms to predict the suitability of Russet potato clones for advancement in breeding trials. Leveraging data from manually collected trials in the state of Oregon, we investigate the potential of a wide variety of state-of-the-art binary classification models. We conduct a comprehensive analysis of the dataset that includes preprocessing, feature engineering, and imputation to address missing values. We focus on several key metrics such as accuracy, F1-score, and Matthews correlation coefficient (MCC) for model evaluation. The top-performing models, namely the multi-layer perceptron (MLP), histogram-based gradient boosting (HGB), and support vector machine (SVM) classifiers demonstrate consistent and significant results. Variable selection further enhances model performance and identifies influential features in predicting trial outcomes. The findings emphasize the potential of machine learning in streamlining the selection process for potato varieties, offering benefits such as increased efficiency, substantial cost savings, and judicious resource utilization. Our study contributes insights into precision agriculture and showcases the relevance of advanced technologies for informed decision-making in breeding programs.
keywords:
Russet potatoes, machine learning, classification, variable selection
@@ 1 Introduction
Potato (Solanum tuberosum L.) is an important food crop in the United Sates with an annual potato production of 392.24 million cwt in 2022. The Pacific Northwest forms the largest sub-region in potato production in the U.S., contributing 60.1% of U.S. potato production according to USDA’s National Agricultural Statistics Service. Potatoes are not only consumed as a fresh vegetable but also form a raw material for the French fry, chip, and starch processing industries. In the U.S., 39% of the potato crop is utilized by the frozen processing sector. The Columbia basin of Oregon and Washington states is the leader in the frozen potato sector, with more than 80% of potatoes grown here being russet type for utilization by the processing industry.
Developing new russet potato varieties with high yield, good processing quality, climate resilience, and disease resistance is the main focus of the Oregon State University’s Potato Breeding and Variety Development program, an integral part of the Northwest Variety Development Program (also known as the Tri-State program). Typically, it takes 12–13 years from the hybridizations to select and release a new potato variety. In the first field year, seedling tubers generated from true potato seeds are planted in Klamath Falls (around 60,000 clones) as single hills and about 1% of the clones selected are planted at Hermiston and Klamath Falls in the second field year. Phenotypic field selections are performed in the first and second years and then the selected clones are tested through Oregon statewide replicated trials at Hermiston, Klamath Falls, and Ontario. The Oregon statewide trials are typically performed in years three to five and the selected clones will then move to a Tri-State trial, and then to regional trials from the sixth or seventh field year to the tenth or eleventh year. Potato varieties released from the Tri-State variety development program such as the “Clearwater Russet”, “Mountain Gem Russet”, “Umatilla Russet”, and “Ranger Russet” have proved to be significantly successful by producing higher total yield and higher percentage of finished product derived from raw tubers in the processing industry. Benefits from the adoption of new russet potato varieties continue to grow over time [1].
@@
The complexity further increases due to the intricate interplay between distinct traits and environmental interactions needed for multi year and multi location trials. Breeders must decide whether the potato clone should go into the subsequent selection trial stage based on its performance or whether retaining a poorly performing clone could add unnecessary cost to the program resulting in decreased breeding selection efficiency. Advances in genetic tools have resulted in the use of molecular markers and genomic selection to improve selection efficiency. Even with molecular and genomic data, it is essential to carry multi year, multi location trials and the breeder must select which clones to advance. In this context, artificial intelligence (AI) or machine learning technologies can help evaluate potato clones and assist in making such decisions [7]. The ascendancy of AI technologies in revolutionizing data analytics is evident, particularly within breeding programs. Researchers have used machine learning models to obtain highly accurate predictions of root yield based on the vegetative indices of the different growth stages of cassava (Manihot esculenta Crantz) [18]. With the potent integration of machine learning algorithms and chemical composition as predictors, researchers have successfully predicted consumer flavor perceptions of blueberries and tomatoes [4]. Data-informed decision making facilitated by machine learning models can assist breeders select potato clones more accurately and reduce the number of varieties proceeding to subsequent stages. The consequential impact includes substantial financial savings, increased breeding selection efficiency, and judicious resource utilization.
@@ 1.1 Related Work
Precision agriculture and the incorporation of advanced technologies like artificial intelligence are becoming crucial for maintaining high-quality yields for many crops. Machine learning methods have been applied to many crops whose information was obtained through various data collection techniques [6, 2].
@@
In a study of French fry potato in Poland, machine learning methods of regression to predict yield using diverse datasets (agronomic, weather, satellite, and soil data) were used [11]. The authors found that consideration of features, coupled with appropriate data preprocessing techniques (e.g., imputation) and outlier management, is crucial for accurate predictions in the field of agricultural yield forecasting.
@@
Another study of potato employed the machine learning models of support vector machines (SVM) and multi-layer perceptrons (MLP) to classify potato varieties based on macroscopic observations and measurements for selection or de-selection during trials much like those used in our study [13]. For each of these two classification techniques, four models were constructed—two for binary classification and two for 4-class classification. The results were reported in terms of the accuracy score (number of correct predictions) with the SVM method demonstrating higher accuracy than MLP in the 4-class models. It is unclear how many features [13] considered or how they addressed missing data. Further, no parameter tuning was performed.
@@ 2.1 Study Area and Data
Figure 1: US State of Oregon. Points represent the cities in which the potato trials were conducted: Corvallis, Hermiston, Ontario, and Klamath Falls. From each of these states, russet potatoes such as the one depicted at the center of the map (source: Wikimedia) were collected and data of each clone’s characteristics were recorded for analysis.
Table 1: Early and late planting season dates over the years.
@@
Year Planting Vine Kill Planting Vine Kill
2013 04/05/13 09/04/13
2014 04/16/14 09/11/14
2015 03/26/15 07/22/15 04/06/15 09/04/15
2016 03/25/16 07/29/16 04/07/16 09/02/16
2017 03/27/17 08/10/17 04/04/17 08/28/17
2018 03/26/18 08/08/18 04/09/18 08/31/18
2019 03/28/19 08/01/19 04/16/19 09/02/19
2020 03/23/20 08/01/20 04/21/20 08/31/20
2021 03/17/21 07/30/21 03/31/21 08/31/21
[…]
@@ 2.2 Feature Engineering
where $T_{\mathrm{MAX}}$ and $T_{\mathrm{MIN}}$ are the day’s maximum and minimum air temperatures, respectively, and $T_{\mathrm{BASE}}$ is “the temperature below which the process of interest does not progress” [12]. In our case, $T_{\mathrm{BASE}}$ is the lowest temperature at which potato continues to grow, or 8 Celsius.
@@ 2.3 Data Imputation
To impute data, we first standardized the dataset and then used Scikit-learn’s [14] function IterativeImputer with sample_posterior=True, which performs multiple imputation by chained equations, or MICE [20]. Under the assumption that the data is missing at random, MICE works by first performing a simple mean imputation to serve as place holders for all variables. For each variable in turn, the place holder means are removed and the variable is treated as dependent on the other independent variables in a regression problem.
@@
Once numerical variables were standardized, we hot encoded the categorical variables - a process where all categorical variables’ values are accounted for all observations, but for an observation’s given category, its value is encoded to 1 while the other non-applicable values are encoded to zero - and handled missing values in a similar manner, i.e., they were either dropped or imputed. We then ran a grid search with cross validation to find the best set of hyperparameters for each of the models considered (see Section 2.5).
@@ 2.5 Classification Models and Evaluation Metrics
    $\boldsymbol{k}$-Nearest neighbors: This is a fundamental and widely applied clustering method in unsupervised learning. The objective of the algorithm is to partition a dataset into $k$ clusters (for a predefined value of $k$) based on similarity, by iteratively computing the Euclidean distance of each data point and cluster representatives, assigning the data points to clusters, and then moving the cluster representative to be the center of the data points it represents. Such an algorithm leads to a local optimum and so it is sensitive to initial conditions and shape assumptions. Also, since clustering relies on computing arithmetic means, this algorithm is sensitive to outliers [17].
@@
    Gaussian process based on Laplace approximation: This approach involves placing a Gaussian Process (GP) prior over a latent function, which is then transformed through the logistic function to obtain a prior on the class probability. The Laplace approximation is employed for posterior inference, approximating the non-Gaussian joint posterior with a Gaussian distribution. With this method, positive training examples will give rise to a positive coefficient for their kernel function, while negative examples will give rise to a negative coefficient, making it analogous to the support vector machine (SVM) model (see Method 10 below). However, all features provide information to perform predictions, making it an inefficient algorithm in high dimensional spaces [21].
@@
    The predicted class variable maximizes the right-hand side of the proportion, given the relative frequency of the class variable in the training dataset [22].
@@
    The decision boundary between each pair of classes $k$ and $l$ is the set of $x$ values satisfying the quadratic equation $\delta_{k}(x)=\delta_{l}(x)$ [9].
@@
    Decision tree: This is a supervised learning approach in which a given vector of explanatory variables $X$ is partitioned into the nodes of a binary tree whose edges represent the mutually exclusive set of outcomes of the parent node. Each edge leads to a different input feature until it reaches the leaf nodes, where a class is assigned to the sequence of feature node outcomes; the goal is to group samples with the same label. Classification depends on the underlying decision tree structure, which is decided by hyperparameters such as the split criteria and the maximum depth of the tree. Although simple to understand, decision trees can be unstable or cause over-fitting depending on the choice of hyperparameters [9].
@@
    Adaptive boosting: This is a commonly used ensemble learning approach in which a sequence of weak learners (models that are slightly better than guessing randomly), such as small decision trees, is repeatedly applied to modified versions of the data (i.e., by changing the weights applied to the training observations). The predictions from all the weak classifiers are then combined to produce a majority vote prediction [9].
@@
    Random forest: This ensemble learning approach is an extension of the bootstrap aggregating (or bagging) method, in which random samples of the training data are selected with replacement and then trained on individually using the base learner. Random forests work similarly by constructing an uncorrelated collection of decision trees. Each tree in the ensemble is constructed from training data sampled with replacement, and an exhaustive search through the features is performed in order to obtain the best split in the node set. Although this is a robust algorithm and might be selected for this reason, it is more complex and time consuming than other ensemble techniques [3].
@@
    Multi-layer perceptron (MLP): This is a feedforward neural network with at least three layers—an input layer, one or more hidden layers, and an output layer—in which neurons are fully connected with a nonlinear activation function. The MLP classifier is trained by backpropagation to optimize a loss function. The Broyden–Fletcher–Goldfarb–Shanno (BFGS) method, the optimization algorithm used in this framework, is an iterative method for solving unconstrained nonlinear optimization problems [8].
@@
    Histogram-based gradient boosting (HGB): Gradient boosting algorithms are additive machine learning models that combine many weak learners such as decision trees to form a strong learner. As finding the best split points for learning the decision trees via sorting the values of each feature is time and memory intensive, the histogram based algorithm buckets continuous feature values into discrete bins and constructs a histogram, thus reducing the number of unique values to sort [10]. Several hyperparameters in the histogram-based gradient boosting classifier (HGBC) can be optimized during learning, such as maximum tree depths, maximum number of bins and the $L_{2}$ regularization parameter, to avoid over-fitting issues while affording good performance.
@@
    If the data is not linearly separable for the desired classification problem, then there will be no hyperplane separating the two classes. However, a nonlinear decision boundary may still exist. As SVM is an inherently linear classification method, we transform the vector space via a kernel trick. A kernel is a function that take as inputs the observations of the original feature space and outputs the dot products of vectors in a transformed target space where finding hyperplanes that separate the data is more likely [16]. Two important hyperparameters in a nonlinear SVM are $C$ and $\gamma$. The scale parameter $\gamma$ determines the radius of influence of a single observation, while $C$ controls the size of slack variables that allow some observations to be outside the target hyperspace. Setting these parameters appropriately is a trade-off between controlling the classification error and overfitting the data.
@@
    An ensemble of classifiers along with a meta-classifier: Also called a stacking classifier, two or more classifiers are applied to the datasets and the resulting classification predictions are used to train a meta-classifier [19]. The stacking classifier used in this study has gradient boosting and SVM as its level-1 classifiers, and logistic regression as its level-2 (or meta-) classifier.

### v3

@@ Abstract
We explore the application of machine learning algorithms specifically to enhance the selection process of Russet potato clones in breeding trials by predicting their suitability for advancement. This study addresses the challenge of efficiently identifying high-yield, disease-resistant, and climate-resilient potato varieties that meet processing industry standards. Leveraging data from manually collected trials in the state of Oregon, we investigate the potential of a wide variety of state-of-the-art binary classification models. The dataset includes 1086 clones, with data on 38 attributes recorded for each clone, focusing on yield, size, appearance, and frying characteristics, with several control varieties planted consistently across four Oregon regions from 2013-2021. We conduct a comprehensive analysis of the dataset that includes preprocessing, feature engineering, and imputation to address missing values. We focus on several key metrics such as accuracy, F1-score, and Matthews correlation coefficient (MCC) for model evaluation. The top-performing models, namely a feedforward neural network classifier (Neural Net), histogram-based gradient boosting classifier (HGBC), and a support vector machine classifier (SVC), demonstrate consistent and significant results. To further validate our findings, we conducted a simulation study using the aims, data-generating mechanisms, estimands, methods, and performance measures (ADEMP) framework, simulating different data-generating scenarios to assess model robustness and performance through true positive, true negative, false positive, and false negative distributions, area under the receiver operating characteristic curve (ROC) and MCC. The simulation results highlight that non-linear models like SVC and HGBC consistently show higher ROC and MCC than logistic regression, thus outperforming the traditional linear model across various distributions, and emphasizing the importance of model selection and tuning in agricultural trials. Variable selection further enhances model performance and identifies influential features in predicting trial outcomes. The findings emphasize the potential of machine learning in streamlining the selection process for potato varieties, offering benefits such as increased efficiency, substantial cost savings, and judicious resource utilization. Our study contributes insights into precision agriculture and showcases the relevance of advanced technologies for informed decision-making in breeding programs.
@@ 1 Introduction
Potato (Solanum tuberosum L.) is an important food crop in the United States with an annual potato production of 392.24 million cwt in 2022. The Pacific Northwest forms the largest sub-region in potato production in the U.S., contributing 60.1% of U.S. potato production according to USDA’s National Agricultural Statistics Service. Potatoes are not only consumed as a fresh vegetable but also form a raw material for the French fry, chip, and starch processing industries. In the U.S., 39% of the potato crop is utilized by the frozen processing sector. The Columbia basin of Oregon and Washington states is the leader in the frozen potato sector, with more than 80% of potatoes grown here being russet type for utilization by the processing industry.
Developing new russet potato varieties with high yield, good processing quality, climate resilience, and disease resistance is the main focus of the Oregon State University’s Potato Breeding and Variety Development program, an integral part of the Northwest Variety Development Program (also known as the Tri-State program). Typically, it takes 12–13 years from the hybridizations to select and release a new potato variety. In the first field year, seedling tubers generated from true potato seeds are planted in Klamath Falls (around 60,000 clones) as single hills and about 1% of the clones selected are planted at Hermiston and Klamath Falls in the second field year. Phenotypic field selections are performed in the first and second years and then the selected clones are tested through Oregon statewide replicated trials at Hermiston, Klamath Falls, and Ontario. The Oregon statewide trials are typically performed in years three to five and the selected clones will then move to a Tri-State trial, and then to regional trials from the sixth or seventh field year to the tenth or eleventh year. Potato varieties released from the Tri-State variety development program such as the “Clearwater Russet”, “Mountain Gem Russet”, “Umatilla Russet”, and “Ranger Russet” have proved to be significantly successful by producing higher total yield and higher percentage of finished product derived from raw tubers in the processing industry. Benefits from the adoption of new russet potato varieties continue to grow over time (Araji and Love, 2002).
@@
The complexity further increases due to the intricate interplay between distinct traits and environmental interactions needed for multi year and multi location trials. Breeders must decide whether the potato clone should go into the subsequent selection trial stage based on its performance or whether retaining a poorly performing clone could add unnecessary cost to the program resulting in decreased breeding selection efficiency. Advances in genetic tools have resulted in the use of molecular markers and genomic selection to improve selection efficiency. Even with molecular and genomic data, it is essential to carry multi year, multi location trials and the breeder must select which clones to advance. In this context, artificial intelligence (AI) or machine learning technologies can help evaluate potato clones and assist in making such decisions (Dwivedi et al., 2021). The ascendancy of AI technologies in revolutionizing data analytics is evident, particularly within breeding programs. Researchers have used machine learning models to obtain highly accurate predictions of root yield based on the vegetative indices of the different growth stages of cassava (Manihot esculenta Crantz) (Selvaraj et al., 2020). With the potent integration of machine learning algorithms and chemical composition as predictors, researchers have successfully predicted consumer flavor perceptions of blueberries and tomatoes (Colantonio et al., 2022).
Data-informed decision making facilitated by machine learning models can assist breeders select potato clones more accurately and reduce the number of varieties proceeding to subsequent stages. The consequential impact includes substantial financial savings, increased breeding selection efficiency, and judicious resource utilization.
@@ 1.1 Related Work
Machine learning has shown promise in revolutionizing the selection process in plant breeding programs. Several studies have demonstrated the potential of integrating advanced data analytics with traditional breeding practices to enhance efficiency and inform decision-making. Machine learning methods have been applied to many crops whose information was obtained through various data collection techniques (Camargo and Smith, 2009; Akbarzadeh et al., 2018).
In a study on cassava (Manihot esculenta Crantz), researchers employed machine learning algorithms to predict root yield based on vegetative indices at different growth stages. The findings highlighted the potential of machine learning in providing insights into the association between above and below-ground traits, offering valuable information for breeding programs (Selvaraj et al., 2020).
Similarly, machine learning methods have been applied to predict consumer flavor perceptions of blueberries and tomatoes. By leveraging chemical composition data and machine learning algorithms, researchers successfully predicted flavor profiles, demonstrating the potential of these techniques in guiding breeding efforts for improved sensory qualities (Colantonio et al., 2022).
@@
In a study of French fry potato in Poland, machine learning methods of regression to predict yield using diverse datasets (agronomic, weather, satellite, and soil data) were used (Kurek et al., 2023). The authors found that consideration of features, coupled with appropriate data preprocessing techniques (e.g., imputation) and outlier management, is crucial for accurate predictions in the field of agricultural yield forecasting.
@@
Another study of potato employed the machine learning models of support vector machines (SVM) and multi-layer perceptrons (MLP) to classify potato varieties based on macroscopic observations and measurements for selection or de-selection during trials much like those used in our study (Ozguven et al., ). For each of these two classification techniques, four models were constructed—two for binary classification and two for 4-class classification. The results were reported in terms of the accuracy score (number of correct predictions) with the SVM method demonstrating higher accuracy than MLP in the 4-class models. It is unclear how many features Ozguven et al. considered or how they addressed missing data. Further, no parameter tuning was performed.
These studies collectively highlight the growing application of machine learning in plant breeding programs. By leveraging advanced data analytics, breeders can gain valuable insights, streamline selection processes, and make informed decisions to accelerate the development of improved crop varieties.
1.2 Contributions of This Study
This study presents a comprehensive evaluation of various classification models for predicting the selection of potato clones in agricultural trials. Our contributions are multifaceted, encompassing both methodological advancements and practical insights:
  1. 1.
    Comprehensive Model Comparison: We systematically compare a diverse set of classification models, including traditional approaches such as Naïve Bayes and decision trees, as well as more advanced techniques like multi-layer perceptron classifiers (MLPC), histogram-based gradient boosting classifiers (HGBC), and support vector machines (SVM). This thorough comparison highlights the strengths and weaknesses of each model within the context of an imbalanced agricultural dataset. By understanding these strengths and weaknesses, we aim to improve the accuracy of selecting potato clones with desirable traits such as yield, disease resistance, and processing quality, which are important for successful breeding outcomes.
  2. 2.
    Evaluation Metrics: By employing the Matthews correlation coefficient (MCC) alongside accuracy and F1-score, we provide a nuanced assessment of model performance. The MCC, in particular, proves to be a valuable metric for evaluating classifiers on imbalanced datasets, offering insights that traditional metrics might overlook. This is particularly important for distinguishing between clones that should be advanced or discarded based on their potential to meet industry standards for yield and quality.
  3. 3.
    Impact of Data Imputation: We investigate the effects of multiple imputation on model performance, providing a detailed analysis of how imputation influences the predictive power of different classifiers. This analysis underscores the importance of handling missing data appropriately in agricultural datasets. Proper imputation allows us to consider all clones participating in the trials, thus allowing classification models to learn from a more diverse dataset and be more generalizable.
  4. 4.
    Forward Feature Selection: Through forward feature selection, we identify key variables that significantly impact model performance. This process not only enhances the predictive accuracy of our top-performing models but also offers valuable insights into the most critical factors influencing potato clone selection.
  5. 5.
    Simulation Study: To validate the robustness of our findings, we conduct a simulation study that replicates various scenarios of data imbalance. This simulation study demonstrates the generalizability of our results and confirms the efficacy of our methodological approach under different conditions.
  6. 6.
    Practical Implications: Our findings have practical implications for potato breeding programs. By identifying the most influential variables and the best-performing models, we provide actionable recommendations that can enhance the efficiency and accuracy of clone selection processes in agricultural trials. Ultimately, this contributes to developing potato varieties that are better suited to meet market demands and environmental challenges, leading to more sustainable agricultural practices.
Overall, this study contributes to agricultural data science by rigorously evaluating classification models to tackle the specific problem of optimizing potato breeding trials. We highlight how selecting appropriate evaluation metrics can improve model performance in identifying clones with desirable traits such as yield, size, appearance, and frying quality. Furthermore, we demonstrate how advanced machine learning techniques can streamline clone selection processes, ultimately reducing costs and improving resource allocation in breeding programs.
@@ 2.1 Study Area and Data
Figure 1: US State of Oregon. Points represent the cities in which the potato trials were conducted: Corvallis, Hermiston, Ontario, and Klamath Falls. From each of these states, russet potatoes such as the one depicted at the center of the map (ZooFari, 2009) were collected and data of each clone’s characteristics were recorded for analysis.
Table 1: Early and late planting season dates over the years. DAP indicates the days after planting for each vine (includes kill date).
@@
Year Planting Vine Kill DAP Planting Vine Kill DAP
2013 04/05/13 09/04/13 152
2014 04/16/14 09/11/14 149
2015 03/26/15 07/22/15 119 04/06/15 09/04/15 152
2016 03/25/16 07/29/16 127 04/07/16 09/02/16 149
2017 03/27/17 08/10/17 137 04/04/17 08/28/17 147
2018 03/26/18 08/08/18 136 04/09/18 08/31/18 145
2019 03/28/19 08/01/19 127 04/16/19 09/02/19 140
2020 03/23/20 08/01/20 132 04/21/20 08/31/20 133
2021 03/17/21 07/30/21 136 03/31/21 08/31/21 154
@@ 2.2 Feature Engineering
where $T_{\mathrm{MAX}}$ and $T_{\mathrm{MIN}}$ are the day’s maximum and minimum air temperatures, respectively, and $T_{\mathrm{BASE}}$ is “the temperature below which the process of interest does not progress” (McMaster and Wilhelm, 1997). In our case, $T_{\mathrm{BASE}}$ is the lowest temperature at which potato continues to grow, or 8 Celsius.
@@ 2.3 Data Imputation
To impute data, we first standardized the dataset and then used Scikit-learn’s (Pedregosa et al., 2011) function IterativeImputer with sample_posterior=True, which performs multiple imputation by chained equations, or MICE (Van Buuren and Groothuis-Oudshoorn, 2011). Under the assumption that the data is missing at random, MICE works by first performing a simple mean imputation to serve as place holders for all variables. For each variable in turn, the place holder means are removed and the variable is treated as dependent on the other independent variables in a regression problem.
Imputation was done after the train/test/validate data splits, ensuring no information leakage between these sets. We validated our imputation technique by running our prediction models on the subset of the original data containing complete information, and on the same subset with values removed randomly following the distribution of missing values for the overall dataset. There was a slight performance improvement with the imputed set, but overall the results between the two sets remained consistent.
@@
Once numerical variables were standardized, we hot encoded the categorical variables—a process where all categorical variables’ values are accounted for all observations, but for an observation’s given category, its value is encoded to 1 while the other non-applicable values are encoded to zero. Currently our framework does not handle imputation on categorical variables, so observation for which categorical values were missing would be dropped. However the only categorical variables in our dataset are “Trial Region” and “Year in Trial”, and there are no missing values for them.
We ran a grid search with cross validation to find the best set of hyperparameters for each of the models considered (see Section 2.5) on all imputed and non-imputed datasets.
@@ 2.5 Classification Models and Evaluation Metrics
    $\boldsymbol{k}$-Nearest neighbors: This is a fundamental and widely applied clustering method in unsupervised learning. The objective of the algorithm is to partition a dataset into $k$ clusters (for a predefined value of $k$) based on similarity, by iteratively computing the Euclidean distance of each data point and cluster representatives, assigning the data points to clusters, and then moving the cluster representative to be the center of the data points it represents. Such an algorithm leads to a local optimum and so it is sensitive to initial conditions and shape assumptions. Also, since clustering relies on computing arithmetic means, this algorithm is sensitive to outliers (Steinbach and Tan, 2009).
@@
    Gaussian process based on Laplace approximation: This approach involves placing a Gaussian Process (GP) prior over a latent function, which is then transformed through the logistic function to obtain a prior on the class probability. The Laplace approximation is employed for posterior inference, approximating the non-Gaussian joint posterior with a Gaussian distribution. With this method, positive training examples will give rise to a positive coefficient for their kernel function, while negative examples will give rise to a negative coefficient, making it analogous to the support vector machine (SVM) model (see Method 10 below). However, all features provide information to perform predictions, making it an inefficient algorithm in high dimensional spaces (Williams and Rasmussen, 2006).
@@
    The predicted class variable maximizes the right-hand side of the proportion, given the relative frequency of the class variable in the training dataset (Zhang, 2004).
@@
    The decision boundary between each pair of classes $k$ and $l$ is the set of $x$ values satisfying the quadratic equation $\delta_{k}(x)=\delta_{l}(x)$ (Hastie et al., 2009).
@@
    Decision tree: This is a supervised learning approach in which a given vector of explanatory variables $X$ is partitioned into the nodes of a binary tree whose edges represent the mutually exclusive set of outcomes of the parent node. Each edge leads to a different input feature until it reaches the leaf nodes, where a class is assigned to the sequence of feature node outcomes; the goal is to group samples with the same label. Classification depends on the underlying decision tree structure, which is decided by hyperparameters such as the split criteria and the maximum depth of the tree. Although simple to understand, decision trees can be unstable or cause over-fitting depending on the choice of hyperparameters (Hastie et al., 2009).
@@
    Adaptive boosting: This is a commonly used ensemble learning approach in which a sequence of weak learners (models that are slightly better than guessing randomly), such as small decision trees, is repeatedly applied to modified versions of the data (i.e., by changing the weights applied to the training observations). The predictions from all the weak classifiers are then combined to produce a majority vote prediction (Hastie et al., 2009).
@@
    Random forest: This ensemble learning approach is an extension of the bootstrap aggregating (or bagging) method, in which random samples of the training data are selected with replacement and then trained on individually using the base learner. Random forests work similarly by constructing an uncorrelated collection of decision trees. Each tree in the ensemble is constructed from training data sampled with replacement, and an exhaustive search through the features is performed in order to obtain the best split in the node set. Although this is a robust algorithm and might be selected for this reason, it is more complex and time consuming than other ensemble techniques (Breiman, 2001).
@@
    Multi-layer perceptron (MLP): This is a feedforward neural network with at least three layers—an input layer, one or more hidden layers, and an output layer—in which neurons are fully connected with a nonlinear activation function. The MLP classifier is trained by backpropagation to optimize a loss function. The Broyden – Fletcher – Goldfarb – Shanno (BFGS) method, the optimization algorithm used in this framework, is an iterative method for solving unconstrained nonlinear optimization problems (Hinton, 1990).
@@
    Histogram-based gradient boosting (HGB): Gradient boosting algorithms are additive machine learning models that combine many weak learners such as decision trees to form a strong learner. As finding the best split points for learning the decision trees via sorting the values of each feature is time and memory intensive, the histogram based algorithm buckets continuous feature values into discrete bins and constructs a histogram, thus reducing the number of unique values to sort (Ke et al., 2017). Several hyperparameters in the histogram-based gradient boosting classifier (HGBC) can be optimized during learning, such as maximum tree depths, maximum number of bins and the $L_{2}$ regularization parameter, to avoid over-fitting issues while affording good performance.
@@
    If the data is not linearly separable for the desired classification problem, then there will be no hyperplane separating the two classes. However, a nonlinear decision boundary may still exist. As SVM is an inherently linear classification method, we transform the vector space via a kernel trick. A kernel is a function that take as inputs the observations of the original feature space and outputs the dot products of vectors in a transformed target space where finding hyperplanes that separate the data is more likely (Schölkopf and Smola, 2001). Two important hyperparameters in a nonlinear SVM are $C$ and $\gamma$. The scale parameter $\gamma$ determines the radius of influence of a single observation, while $C$ controls the size of slack variables that allow some observations to be outside the target hyperspace. Setting these parameters appropriately is a trade-off between controlling the classification error and overfitting the data.
@@
    An ensemble of classifiers along with a meta-classifier: Also called a stacking classifier, two or more classifiers are applied to the datasets and the resulting classification predictions are used to train a meta-classifier (Tang et al., 2014). The stacking classifier used in this study has gradient boosting and SVM as its level-1 classifiers, and logistic regression as its level-2 (or meta-) classifier.